OpenStudy (anonymous):

**PLEASE HELP! FAN AND MEDAL!!!** 4/5x + 4/3 = 2x solve for X

OpenStudy (solomonzelman):

4/5x + 4/3 = 2x bring 4/5x to the left side first

OpenStudy (anonymous):

x= 10/9

OpenStudy (solomonzelman):

lovatic4life, don't just give the answer!

OpenStudy (anonymous):

hes my budy ill do what i want

OpenStudy (anonymous):

so how do I bring 4/5 to the left side?

OpenStudy (anonymous):

*4/5x

OpenStudy (solomonzelman):

4/5x to the left side. and you subtract it from both sides basically.

OpenStudy (solomonzelman):

4/5x + 4/3 -4/5x = 2x -4/5x OK?

OpenStudy (anonymous):

ok so then I have 4/3 = 2x - 4/5x?

OpenStudy (solomonzelman):

Yes, and then 2x-4/5x is what?

OpenStudy (anonymous):

I have no idea

OpenStudy (solomonzelman):

2x-4/5x OR